OpenStudy (anonymous):

what are the solutions of the equation 2(x-2)+2(x-4)=2x?

OpenStudy (anonymous):

evaluate the brackets first

OpenStudy (anonymous):

how do I solve it step by step?

OpenStudy (anonymous):

2(x-2)+2(x-4)=2x evaluating brackets we get 2x - 4 + 2x -8 = 2x 2x+2x -4 - 8 = 2x 4x -12 = 2x 4x-2x = 12 2x = 12 x = 6

OpenStudy (anonymous):

ok I got that but what's the other solution too? Or is it 0?

OpenStudy (anonymous):

its a linear equation it has only one solution